OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83?

OpenStudy (cwrw238):

use 35th term = 13 + 34d d is found by (a17 - a1) / 16 = (-83 - 13) / 16

OpenStudy (cwrw238):

the terms of an AS are a, a+d, a + 2d, a+ 3d ......

OpenStudy (anonymous):

theres no way its -6 but thats what i got when i did (-83-13)/16

OpenStudy (cwrw238):

d = -6 plug this into formula in my first post

OpenStudy (cwrw238):

= 13 + 34(-6)

OpenStudy (anonymous):

oh d = -6 .. okay. so then, 13 + 34(-6) ?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

okay hold on. i dont have a calculator :/

OpenStudy (anonymous):

-191 ?

OpenStudy (cwrw238):

13 - 204 = 191 is right

OpenStudy (anonymous):

sweet. thanks!

OpenStudy (cwrw238):

-191 i mean

OpenStudy (cwrw238):

yw